An empty railroad car traveling at a speed of 21 m
Answer #1

m1v1 + m2v2 = (m1+m2)v

v = (21m1 + 0m2)/(m1 + m2)

v = (21m1)/3m1 {m2 = 2m1}

v = 7 m/s